OpenStudy (anonymous):

This is a monster... Cal 3 if anyone is interested. I have to convert a triple integral from rectangular to cylindrical bounds and evaluate. Integral in the replies.

OpenStudy (anonymous):

\[\int\limits_{-1}^{1} \int\limits_{0}^{\sqrt{1-y}} \int\limits_{0}^{x} x^2+y^2 dz dx dy\]

OpenStudy (anonymous):

\[\begin{cases}x=r\cos\theta\\y=r\sin\theta\\ x^2+y^2=r^2\\z=z\\ dz~dx~dy=r~dr~d\theta\end{cases}\]